# Provably Mitigating Overoptimization in RLHF: Your SFT Loss is Implicitly an Adversarial Regularizer

### v2

@@ 1 Introduction
Figure 1: Left: Reward overoptimization due to the distributional shift and uncertainty in reward. Right: Overoptimization can cause the probability of outputting the preferred responses from preference data to decrease substantially using original DPO proposed by (Rafailov et al., 2023). Our algorithm (RPO) significantly alleviates this decrease. See more discussions in Section 6.
@@ 1.2 Related Works
Finally, we mention that the algorithm design of RPO is also related to the “pessimism” principle in the standard offline RL literature. It proposes to maintain a pessimistic estimate of the policy values or constrain the policy not to take unseen actions in the data to handle the challenge of the insufficient coverage of the dataset, e.g., (Jin et al., 2021; Uehara and Sun, 2021; Xie et al., 2021a, b; Rashidinejad et al., 2021; Yin and Wang, 2021; Xiong et al., 2022; Liu et al., 2022a; Shi et al., 2022; Zhan et al., 2022; Lu et al., 2022; Rashidinejad et al., 2022; Shi and Chi, 2024; Blanchet et al., 2024; Liu et al., 2022b; Liao et al., 2021; Fu et al., 2022). In contrast, we consider the offline RLHF problem and the techniques to obtain the objective of the RPO algorithm (see Section 4) along with its sample complexity analysis are new and different from these prior works on standard offline RL.
@@ 5.1 Establishing the Sample Complexity of Maximin Objective (Algorithm 1)
Taking the policy class $\Pi$ as (2.4), supposing that Assumptions 5.1 and 5.2 hold, and assuming that the reward model class $\mathcal{R}$ has a finite $\varepsilon$-epsilon covering number under $\|\cdot\|_{\infty}$-norm $\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})<+\infty$ with $\varepsilon=(6\cdot(1+e^{R})\cdot N)^{-1}$. Setting
@@ Appendix B Proofs for Sample Complexity Analysis
B.1 Proof of Theorem 5.3
@@
$\qquad=\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x),a^{0}\sim\pi^{\mathrm{ref}}(\cdot|x)}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (B.3)
@@
$\qquad=\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{ref}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\}$ (B.21)

### v3

@@ 1 Introduction
action $a$ $b$ $c$
$r^{\star}$ 1 0.5 0
Dataset $\mathcal{D}=\{(a,b,1)\}$
$\pi^{\mathrm{ref}}$ $0.45$ $0.45$ $0.1$
$\pi^{\mathrm{DPO}}$ $0.5$ 00 0.5
$\pi^{\mathrm{RPO}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bf 1.0}$ 00 00
Figure 1: Left: Reward overoptimization due to distributional shift and uncertainty in reward. Middle: Overoptimization can cause the probability of outputting the preferred responses from preference data to decrease substantially using original DPO proposed by (Rafailov et al., 2023). Our algorithm (RPO) can significantly alleviate this decrease. See more discussions in Section 6. Right: For the illustration of the intuition behind the success of RPO, we give a simple example with no prompts and only three responses $a$, $b$, and $c$ (Xu et al., 2024). The dataset $\mathcal{D}$ only consists of a single data point where response $a$ is preferred than response $b$ (as indicated by the label $1$). Writing out the DPO loss, one needs to minimize $\log(1+(\pi(b)/\pi(a))^{\beta})$ for $0\leq\pi(a)+\pi(b)\leq 1$, which gives $\pi(b)=0$ and leaves $\pi(a)$ arbitrary. Thus possible solutions to DPO objective would bring out non-zero probabilities of response $c$ that is not desirable since $r^{\star}(c)=0$. In contrast RPO objective simultaneously maximizes another term $\log\pi(a)$ as regularizer, avoiding the output of response $c$.
@@ 1.2 Related Works
Finally, we mention that the algorithm design of RPO is also related to the “pessimism” principle in the standard offline RL literature. It proposes to maintain a pessimistic estimate of the policy values or constrain the policy not to take unseen actions in the data to handle the challenge of the insufficient coverage of the dataset, e.g., Jin et al. (2021); Uehara and Sun (2021); Xie et al. (2021a, b); Rashidinejad et al. (2021); Yin and Wang (2021); Fu et al. (2022); Xiong et al. (2022); Liu et al. (2022a, b); Shi et al. (2022); Zhan et al. (2022); Lu et al. (2022); Rashidinejad et al. (2022); Shi and Chi (2024); Blanchet et al. (2024). In contrast, we consider the offline RLHF problem and the techniques to obtain the objective of our algorithm along with its sample complexity analysis are new and different from these prior works on standard offline RL.
@@ 5.1 Establishing the Sample Complexity of Maximin Objective (Algorithm 1)
Taking the policy class $\Pi$ as (2.4), supposing that Assumptions 5.1 and 5.2 hold, and assuming that the reward model class $\mathcal{R}$ has a finite $\varepsilon$-covering number under $\|\cdot\|_{\infty}$-norm $\mathcal{N}_{\varepsilon}(\mathcal{R},\|\cdot\|_{\infty})<+\infty$ with $\varepsilon=(6\cdot(1+e^{R})\cdot N)^{-1}$. Setting
@@ Appendix B Proofs for Sample Complexity Analysis
B.1 Concepts
We provide some useful concepts for the simplicity of later discussions.
  • Hellinger distance $D_{\mathrm{Hellinger}}(p\|q)$ between two probability density functions $p$ and $q$ defined on $\mathcal{X}$ is defined as
    $D_{\mathrm{Hellinger}}(p\|q)=\frac{1}{2}{\int_{x\in\mathcal{X}}\left(\sqrt{p(x)}-\sqrt{q(x)}\right)^{2}}{}\mathrm{d}x.$
  • Total variation (TV) distance $D_{\mathrm{TV}}(p\|q)$ between two probability density functions $p$ and $q$ defined on $\mathcal{X}$ is defined as
    $D_{\mathrm{TV}}(p\|q)=\frac{1}{2}{\int_{x\in\mathcal{X}}|p(x)-q(x)}|\mathrm{d}x.$
  • Kullback–Leibler (KL) divergence $\text{KL}(p\|q)$ between two probability density functions $p$ and $q$ defined on $\mathcal{X}$ is defined as
    $\text{KL}(p\|q)=\int_{x\in\mathcal{X}}\log\left(\frac{p(x)}{q(x)}\right)p(x)\mathrm{d}x.$
  • We denote $\mathcal{N}_{\epsilon}(\mathcal{F},\|\cdot\|_{\infty})$ as the $\epsilon$-covering number (Zhou, 2002) for function class $\mathcal{F}$ under the infinity norm $\|\cdot\|_{\infty}$. Widely used in the theoretical analysis, the $\epsilon$-covering number characterizes the complexity of the function class $\mathcal{F}$.
B.2 Proof of Theorem 5.3
@@ B.1 Concepts
$\qquad=\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\}$ (B.21)